Naina runs a business that only sells hamburgers in a building she owns. Every month, they sell 6,000 hamburgers at $5 per hamburger. She spends $2 per hamburger on supplies (meat, lettuce, bread, etc.). She also pays John and Rajeev each $2500 per month to work at the restaurant. Finally, utilities cost $500 per month. Susana works full-time at the restaurant and keeps the accounting profits for herself. What is the accounting profit of the business? Susana could rent out her building for $5000 per month. She can also make $6000 per month as an accountant. Based on this, what is the economic profit of her business?


1
Expert's answer
2021-05-03T10:53:36-0400

The accounting profit of the business:


"\pi=TR-TC"

"TR=p\times Q =6000\times5=30000"

"TC=FC+VC"

"FC=2\times2500+500=5500"

"VC=AVC\times Q=2\times6000=12000"

"TC=5500+12000=17500"

"\pi=30000-17500=12500"

The economic profit:


"12500-(6000+5000)=1500"